# AI-Native 6G Physical Layer with Cross-Module Optimization and Cooperative Control Agents

### v1

@@ VI Conclusion
This article proposes a framework of AI-native cross-module optimized physical layer with cooperative control agents, encompassing holistic optimization across physical layer AI/ML modules integrated with multiple enhancement mechanisms and control strategies. Specifically, simultaneous optimization across global modules breaks down traditional inter-module information silos to facilitate end-to-end training toward global objectives. Moreover, AI/ML-based cross-layer modulation breaks the isolation of inter-layer features, thus expanding the constellation solution space. The theoretical analysis that reveal the rationality of cross-layer modulation is also provided. The utility-oriented precoder construction directly generates the precoder optimized for end-to-end performance. By integrating modulation into AI/ML-enabled CSI feedback, the bit-level bottlenecks are effectively bypassed. Additionally, AI/ML-based control agents facilitate global throughput optimization with integrated control mechanisms. Finally, simulation results using practical 3GPP assumptions demonstrate the superiority of the proposed solutions in terms of block error rate and throughput. Above contents of this article also hopefully provide valuable insights for 3GPP discussions in the future. In the future work, to maintain scalability, several approaches expanded from on our proposed framework can be considered, such as subband-level model processing partial bandwidth segments sequentially and pretraining a base model offline followed by efficient online fine-tuning with limited samples, etc. We also recognize the importance of comprehensive channel estimation robustness in practical deployments, and are actively investigating AI-native solutions for more practical scenarios including ICI and ISI mitigation as a direction for further research.

### v2

@@ VI Conclusion
This article proposes a framework of AI-native cross-module optimized physical layer with cooperative control agents, encompassing holistic optimization across physical layer AI/ML modules integrated with multiple enhancement mechanisms and control strategies. Specifically, simultaneous optimization across global modules breaks down traditional inter-module information silos to facilitate end-to-end training toward global objectives. Moreover, AI/ML-based cross-layer modulation breaks the isolation of inter-layer features, thus expanding the constellation solution space. The theoretical analysis that reveals the rationality of cross-layer modulation is also provided. The utility-oriented precoder construction directly generates the precoder optimized for end-to-end performance. By integrating modulation into AI/ML-enabled CSI feedback, the bit-level bottlenecks are effectively bypassed. Additionally, AI/ML-based control agents facilitate global throughput optimization with integrated control mechanisms. Finally, simulation results using practical 3GPP assumptions demonstrate the superiority of the proposed solutions in terms of block error rate and throughput. Above contents of this article also hopefully provide valuable insights for 3GPP discussions in the future. In the future work, to maintain scalability, several approaches expanded from our proposed framework can be considered, such as subband-level model processing partial bandwidth segments sequentially and pretraining a base model offline followed by efficient online fine-tuning with limited samples, etc. We also recognize the importance of comprehensive channel estimation robustness in practical deployments, and are actively investigating AI-native solutions for more practical scenarios including ICI and ISI mitigation as a direction for further research.